Let V(x) denote the number of litres of fuel left in an aircraft’s fuel tank if it has flown x km. Suppose that V(x) satisfies the following differential equation: V (x) = −aV(x) − b. Here, the fuel consumption per km is a constant b > 0. The term −aV(x), with a > 0, is due to the weight of the fuel) a) solve the equation with v(0)=vo b) how many km,x,can the plane fly if it takes off with vo litres in the tank


1
Expert's answer
2022-01-18T11:47:37-0500

a)

"V '(x) = \u2212aV(x) \u2212 b"


"\int \frac{dV}{aV+b}=-x+c"


"ln(aV+b)\/a=-x+c"


"ln(aV_0+b)\/a=c"


"ln(aV+b)\/a=-x+ln(aV_0+b)\/a"


"x=\frac{ln\frac{aV_0+b}{aV+b}}{a}"


b)

for "V=0" :


"x=\frac{ln\frac{aV_0+b}{b}}{a}"